OpenStudy (anonymous):

*Please help* Express each ratio as a fraction in the simplest form. 1) 4 quarts 14 quarts 2) 7 dimes to 84 dimes 3) 10 miles out of 20 miles 4) 8 blue cars out of 44 cars

OpenStudy (mertsj):

Do you know what a ratio is?

OpenStudy (anonymous):

yes i do

OpenStudy (mertsj):

Do you know how to reduce fractions?

OpenStudy (anonymous):

no i just learned it. Im not so sure.

OpenStudy (mertsj):

Write the first ratio in fraction form: \[\frac{4}{14}\]

OpenStudy (mertsj):

Can you think of a number that divides into both 4 and 14?

OpenStudy (anonymous):

um 2

OpenStudy (mertsj):

\[\frac{4}{14}=\frac{2\times 4}{2 \times 7}\]

OpenStudy (mertsj):

Now, of course, 2/2=1 so cancel the 2's and you're done.

OpenStudy (anonymous):

so 4/7 ?

OpenStudy (mertsj):

\[\frac{4}{14}=\frac{2 \times 2}{2 \times 7}\]

OpenStudy (mertsj):

Sorry for the typo...2/7

OpenStudy (anonymous):

ah no its fine. I get it now. Thanks for the help!!

OpenStudy (mertsj):

yw